11. [-13.22 Points] DETAILS LARLINALG8 2.6.017. Consider the following. 5 (1,5) 4 (2, 4) 3 2 (2, 2) (3, 1) - Х - 1 2 3 4 -14(a) Sketch the line that appears to be the best fit for the given points.

(b) Find the least squares regression line. y(x)=

(c) Calculate the sum of squared error.
Answer #1

(a)

The figure below illustrates a scatter plot and a line of best fit.

The equation of the line is, y = mx + c , where m is the slope of the line, and c is the y-intercept.

The line passes through the points, A(1,5), and, B(3,1).

Slope, m=\frac{5-1}{1-3}

i.e., m=-2

Therefore the equation of the line becomes, y=-2x+c

The above line passes through the point, A(1,5). Using the co-ordinates of the point A(1,5), in the above, equation of the line, we have,

5=(-2)(1)+c

i.e., c=7

Substituting the value of, c=7 in the above, equation of the line, we have, the equation of the line,

y=-2x+7

The equation of the line of best fit, is, y=-2x+7 .

7 CO -5 +3 N -4 -3 -2 -1 0 2 3 4 5 6 00 9 10 1

(b)

The least squares regression line is given by the formula,

\hat{y}=b_0+b_1x,

Its slope bi and y-intercept b_0 are computed using the formulas, b_1=\frac{SS_{xy}}{SS_{xx}} , and, b_0=\bar{y}-b_1\bar{x} , where,

SS_{xx}=\sum x^2-\frac{1}{n}\left ( \sum x \right )^2, SS_{xy}=\sum xy-\frac{1}{n}\left ( \sum x \right )\left ( \sum y \right )

T is the mean of all the x-values, \bar{y} is the mean of all the y-values, and n is the number of pairs in the data set.

x y x^2 xy
1 5 1 5
2 2 4 4
2 4 4 8
3 1 9 3
\sum 8 12 18 20

In the last line of the table above, we have the sum of the numbers in each column. Using them we compute:

SS_{xx}=\sum x^2-\frac{1}{n}\left ( \sum x \right )^2=18-\frac{1}{4}(8)^2=2

SS_{xy}=\sum xy-\frac{1}{n}\left ( \sum x \right )\left ( \sum y \right )=20-\frac{1}{4}(8)(12)=-4

\bar{x}=\frac{\sum x}{n}=\frac{8}{4}=2

\bar{y}=\frac{\sum y}{n}=\frac{12}{4}=3

and we now have,

b_1=\frac{SS_{xy}}{SS_{xx}}=\frac{-4}{2}=-2

Using the value of b_1=-2 , \bar{x}=2 , and, \bar{y}=3 in the equation of the least squares regression line, we have, b_0=\bar{y}-b_1\bar{x}=3-(-2)(2)=7

We now have the least squares regression line for given data and it is, \hat{y}=7-2x

(c)

We now calculate the sum of squared error.

\textup{SSE}=\sum_{i=1}^{n}(x_i-\bar{x})^2

T y x_i-\bar{x}=x_i-2 (x_i-\bar{x})^2
1 5 -1 1
2 2 0 0
2 4 0 0
3 1 1 1
\sum 8 12 0 2

\textup{SSE}=\frac{(x_1-2)^2+(x_2-2)^2+(x_3-2)^2+(x_4-2)^2}{4}

i.e., \textup{SSE}=\frac{(1-2)^2+(2-2)^2+(2-2)^2+(3-2)^2}{4}

i.e., \textup{SSE}=\frac{1+0+0+1}{4}

i.e., \textup{SSE}=\frac{2}{4}=\frac{1}{2}